OpenStudy (anonymous):

Kendra has $7.05 in nickels and quarters. If she has 15 more quarters than nickels, how many of each coin does she have?

Mehek (mehek14):

15 quarters = $3.75 now add a nickel each time remember every time you add a nickel, add a quarter

OpenStudy (anonymous):

@Mehek14 Is there an equation to solve this?

Mehek (mehek14):

no idrk

Mehek (mehek14):

3.75 + .25 +.05 = 4.05 16 quarters and 1 nickel 4.05 + .25 + .05 =4.35 17 quarters and 2 nickels 4.35 + .25 + 0.05 = 4.65 18 quarters and 3 nickels 4.65 + 0.25 + 0.05 = 4.95 19 quarters and 4 nickels 4.95 + 0.25 + 0.05 = 5.25 20 quarters and 5 nickels 5.25 + 0.25 + 0.05 = 5.55 21 quarters and 6 nickels 5.55 + 0.25 + 0.05 = 5.85 22 quarters and 7 nickels 5.85 + 0.25 + 0.05 = 6.15 23 quarters and 8 nickels 6.15 + 0.25 + 0.05 = 6.45 24 quarters and 9 nickels 6.45 + 0.25 + 0.05 = 6.75 25 quarters and 10 nickles 6.75 + 0.25 + 0.05 = 7.05 26 quarters and 11 nickles